What is the NPV of a project that has an initial investment of $100,000, and expected cash flows of $45,000 each year for years 1 through 3? Use an 8% hurdle (aka discount) rate. Is this a project you should accept or reject? Calculate the Profitability Index for the project as well.
Answer #1

DZ f oc =SUM(D3:06) B 3 2 year Cash flows pv@8% 0 $ (1,00,000.00) 4 1 $ 45,000.00 5 2 $ 45,000.00 6 3 $ 45,000.00 Present val

fx =SUM(D3:06) D7 А D 1 pv@8% 2 year 30 41 5 2 63 Cash flows -100000 45000 45000 45000 =C3/1.08 =C4/1.08 =C5/1.08 NPV Present